OpenStudy (anonymous):

I was wondering how to find these things about the Singapore Flyer (Ferris Wheel)? - Measure of a central angle in degrees - Measure of a central angle in radians - Arc length between two cars or compartments - Area of a sector between two cars or compartments

OpenStudy (anonymous):

Please, I need help :D

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

No

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well I'd expect you would need an image and a protractor... measure the angle... in degrees and then convert to radians.

OpenStudy (anonymous):

for this assignment there was no picture given but I have the diameter, circumference, area of the wheel, and number of compartments on the ferris wheel. @campbell_st

OpenStudy (campbell_st):

ok... so how many compartments...?

OpenStudy (anonymous):

28, diameter is 150 meters, circumference = 471, area of wheel = 17662.5

OpenStudy (anonymous):

28 compartments

OpenStudy (campbell_st):

ok... so a circle contains 360 degree so the angle between compartments is 360/28 does that make sense

OpenStudy (anonymous):

Yes

OpenStudy (campbell_st):

so when you get the sector angle it can be used to get the fraction of the circle sector area = (sector angle)/360 * area arc length = (sector angle)/360 * circumference angle in radians use \[180^o = \pi\] so then \[1^o = \frac{\pi}{180}\] so the measure in radians will be \[(sector \angle) \times \frac{\pi}{180}\] hope it helps

OpenStudy (anonymous):

so 12.86 is the sector angle , then the sector area = 630.80, arc length = 5.36, then I got lost :( @campbell_st @amistre64

OpenStudy (campbell_st):

ok... so the conversion to radians just do \[12.86 \times \frac{\pi}{180} = \]

OpenStudy (anonymous):

Okay so just to clarify the central angle is 12.86 and in radians it is .22? Thank you so much btw, it really has helped loads!! @campbell_st and sorry if I am annoying the crap out of you :P

OpenStudy (campbell_st):

thats ok.... glad to help

OpenStudy (campbell_st):

yes your radian measure is correct